OpenStudy (anonymous):

which is a zero of f(x)=x^2-15=54

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (dumbcow):

factor

OpenStudy (anonymous):

\[x^2-15=54\] \[x^2=69\] \[=\pm \sqrt{69}\]

OpenStudy (dumbcow):

or that, i thought there was an x there haha